Suppose that the short-run world demand and supply elasticities for crude oil are -0.076 and 0.088, respectively. The current price per barrel is $30 and the short-run equilibrium quantity is 23.84 billion barrels per year. Derive the linear demand and supply equations. Also verify your answer



1
Expert's answer
2021-09-13T11:20:42-0400

Linear demand equation is: "Q = a - bP"

Elasticity of demand "= \frac{dQ}{dP} \times \frac{P}{Q} = -b \times \frac{P}{Q}"

When

"P = 30 \\

Q = 23.84

-0.076 = -b \times \frac{30}{23.84} \\

b = 0.06"

From demand function,

"23.84 = a - (0.06 \times 30) \\

23.84 = a - 1.8 \\

a = 25.64"

Linear demand equation: "Q = 25.64 - 0.06P"


Linear supply equation: "Q = c + dP"

Elasticity of supply "= \frac{dQ}{dP} \times \frac{P}{Q} = d \times \frac{P}{Q}"

When

"P = 30 \\

Q = 23.84 \\

0.088 = d \times \frac{30}{23.84} \\

b = 0.07"

From supply function,

"23.84 = c + (0.07 \times 30) \\

23.84 = c + 2.1 \\

c = 21.74"

Linear demand equation: "Q = 21.74 + 0.07P"